I am having a problem finding a line that goes through (4,6)(5,8)
I am using y2-y2 over x2-x1
8-6= 2
5-4= 1
Then y=mx+b
y-5= 5(x+b) I get lost from here unless I am missing something. Any help would be appreciate. Thanks in advnace.

Answer by scott8148(5891)

you are close

you should be using point-slope but you have slope-intercept in there also

the slope is 2 (good work), so just take one of the points and write the equation

y-(y coordinate)=(slope)(x-(x coordinate)) ... y-8=2(x-5)